OpenStudy (anonymous):

let c = log base3 10 and d = log base3 5. express the following in terms of c and d. log base log base 27 50.

OpenStudy (anonymous):

i am understanding a lot more today azteck

OpenStudy (anonymous):

\[c=\log_{3} 10\] \[3^c=10\] \[d=\log_{3} 5\] \[3^d=5\] \[\large \log_{\frac{1}{3}} 5=\log_{3^{-1}} 5\] \[3^{-1}=5\] Therefore that is basically equal to d. \[\large \log_{\frac{1}{3}} 5=d\]

OpenStudy (anonymous):

I reckon this was simpler than the previous one you gave me...

OpenStudy (anonymous):

This was just using simple indices. A third is just equal to 3^-1

OpenStudy (anonymous):

Let me double check this first.

OpenStudy (anonymous):

Now you corrected the question....

OpenStudy (anonymous):

log base log base 27 50?

OpenStudy (anonymous):

@dawnd4fun

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

log base log base? or is it just log base 27 50?

OpenStudy (anonymous):

it is log base27 50

OpenStudy (anonymous):

same c and d

OpenStudy (anonymous):

Is this a different question?

OpenStudy (anonymous):

i did logbase9 10 and got a good answer

OpenStudy (anonymous):

i also have log base9 1/1000

OpenStudy (anonymous):

yes it is different questions, but same c and d parts

OpenStudy (raden):